# MobileDiffusion: Subsecond Text-to-Image Generation on Mobile Devices

### v1

@@ Abstract
The deployment of large-scale text-to-image diffusion models on mobile devices is impeded by their substantial model size and slow inference speed. In this paper, we propose MobileDiffusion, a highly efficient text-to-image diffusion model obtained through extensive optimizations in both architecture and sampling techniques. We conduct a comprehensive examination of model architecture design to reduce redundancy, enhance computational efficiency, and minimize model’s parameter count, while preserving image generation quality. Additionally, we employ distillation and diffusion-GAN finetuning techniques on MobileDiffusion to achieve 8-step and 1-step inference respectively. Empirical studies, conducted both quantitatively and qualitatively, demonstrate the effectiveness of our proposed techniques. MobileDiffusion achieves a remarkable sub-second inference speed for generating a $512\times 512$ image on mobile devices, establishing a new state of the art.
Figure 1: MobileDiffusion text-to-image generation in (a) 8 steps (cfg-aware distillation [23]), and (b) 1 step (UFOGen finetuning [57]).
@@ 1 Introduction
Text-to-image diffusion models  [34, 38, 42, 1, 39, 36] have exceptional capabilities in generating high-quality images conditioned on texts. These models serve as the foundation for a variety of applications, including image editing [12, 50], controlled generation  [60, 33], personalized content generation [41, 9], video synthesis [3, 10], and low-level vision tasks [55, 22]. These large-scale models are essentially considered for necessarily running on servers with powerful neural compute units. Only a few work [23, 20] has barely touched designing diffusion models for mobile devices, which remains a open challenge.
@@
The inefficiency of text-to-image diffusion models arises from two primary factors. Firstly, the inherent design of diffusion models requires iterative denoising to generate images, necessitating multiple evaluations of the model  [14, 46]. Secondly, the complexity of the network architecture in text-to-image diffusion models involves a substantial number of parameters, often reaching into the billions, resulting in computationally expensive evaluations. These combined factors contribute to the high cost associated with deploying text-to-image diffusion models, limiting their practical utility in real-world scenarios. Moreover, the inefficiency challenge poses a significant barrier to the deployment of these models in resource-constrained environments, such as on mobile devices. As a result, despite the potential benefits of deploying generative models on mobile devices, such as enhancing user experience and addressing emerging privacy concerns, this aspect remains relatively unexplored within the current literature.
The optimization of inference efficiency in text-to-image diffusion models has been a active research area. Previous studies predominantly concentrate on addressing the first factor, seeking to reduce the number of network evaluations. Leveraging advanced numerical solvers [28, 29, 2, 19, 45] or distillation techniques  [43, 32, 47, 23], the necessary sampling steps have significantly reduced from several hundreds to around 10. Some recent techniques  [30, 25, 57] even push the boundaries to a few or a single step. Despite these advancements, the deployment of diffusion models on edge devices is still a substantial challenge. On mobile devices, even a minimal number of evaluation steps can be slow or impractical, due to the second factor elucidated earlier — the computational intricacies associated with the model’s architecture. Unfortunately, the architectural efficiency of text-to-image diffusion models has received comparatively less attention. A handful of earlier works briefly touch upon this matter, involving the removal of redundant neural network blocks  [23, 20] or the reorganization of blocks [15] for enhanced efficiency. However, these efforts lack a comprehensive analysis of each component within the model architecture, thereby falling short of providing a holistic guide for designing highly efficient architectures.
Effectively overcoming the challenges imposed by limited computational power of on-device requires an in-depth and holistic exploration of the model’s architectural efficiency. In pursuit of this objective, our research undertakes a detailed examination of each constituent and computational operation within the widely utilized UNet architecture [40]. Consequently, we present a comprehensive guide for crafting highly efficient text-to-image diffusion models. Drawing on these insights, our work introduces MobileDiffusion, embodying a delicately designed model architecture with fewer than 400 million parameters. This lightweight model emerges as a compelling choice for on-device applications, excelling not only in text-to-image generation but also in diverse downstream tasks. Furthermore, the resulting model seamlessly integrates with the techniques for reducing sampling steps discussed earlier. Notably, we successfully applied the sampling techniques [23, 57] to our model. The synergy of efficient architecture design and sampling step reduction empowers us to generate high-quality $512\times 512$ images in sub-second on mobile devices, exemplified by 0.2 second on iPhone 15 Pro. This remarkable result significantly surpasses the previous state-of-the-art in on-device text-to-image generation [23].
@@ 2.1 Architecture Efficiency
A limited number of prior works have discussed the architectural efficiency of diffusion models. [15] proposed the UViT architecture for training diffusion models on high-resolution images. They demonstrated that superior quality and efficiency could be achieved by scaling transformer blocks at the low-resolution segment of the UNet. [20] introduced an approach to distill larger diffusion models into smaller student models by selectively removing specific blocks from the teacher model. Meanwhile, [23] presented an efficient architecture search method. They trained a UNet with redundant blocks using robust training techniques [17, 58] and then pruned certain blocks based on metrics, resulting in an architecture suitable for distillation. In our work, we align with the insights of [15] regarding the transformer components in the UNet while introducing distinctive perspectives elaborated in Section 3. In contrast to approaches like [23, 20], our focus extends beyond mere block removal. Instead, we conduct a more nuanced analysis and modification of the architecture. Additionally, unlike [23, 20], we opt not to employ knowledge distillation from a larger model. Our observations indicate that training our model from scratch yields satisfactory results.
Our research is also closely related the broader domain of efficient architecture design, such as MobileNets [16], ConvNeXt [26], and Efficient Transformers [49].
2.2 Sample Efficiency
A fundamental challenge linked to diffusion models is the slow inference brought by the iterative denoising process. In response to this challenge, diverse strategies have emerged to expedite the required sampling speed, broadly falling into two categories. The first involves developing fast solvers to more efficiently solve the differential equation tied to the denoising process, thereby reducing the necessary discretization steps steps [28, 28, 19, 2, 8]. The second approach leverages knowledge distillation techniques to compress the sampling trajectory  [31, 43, 47, 23]. Empirically, these techniques have demonstrated the ability to decrease the number of sampling steps to around 10 for latent diffusion models like Stable Diffusion (SD) [39]. However, further reductions often lead to a substantial drop in quality. Recent advancements, as presented in works like [30, 25, 57], introduce novel technologies pushing the boundaries of sampling step reduction to a few or even single steps.
3 Mobile Diffusion
In this section, we present our recipe for crafting highly efficient text-to-image diffusion models, which ultimately lead to sub-second generation on mobile device. We start with a brief recap of Diffusion Models in Section 3.1. In section 3.2, we delve into a thorough investigation of UNet architecture optimization, with the goal of enhancing both parameter and runtime efficiency, while ensuring high-quality generation. In Section 3.3, we summarize the candidate architectures of Mobile Diffusion. Finally in Section 3.4, drawing inspiration from recent advancements in diffusion model distillation and diffusion GAN hybrids, we elucidate the methods we adopted for improving sample efficiency.
3.1 Preliminary: Diffusion Models
Diffusion models are trained to gradually transform a Gaussian noise to a sample from target data distribution $p(x)$ via a learned denoising process. After some derivation [14, 46], the training can be interpreted as a denoising process. Specifically, given a data point $x_{0}\sim p(x)$, we can follow a schedule to add Gaussian noise $\epsilon$ to corrupt the clean data into a noisy version $x_{t}$, and the denoising network $\epsilon_{\theta}(x_{t},t)$ is trained to predict the additive noise $\epsilon$
$\mathcal{L}=\mathbb{E}_{x,t,\epsilon\sim\mathcal{N}(0,\mathbf{I})}||\epsilon-\epsilon_{\theta}(x_{t},t)||_{2}^{2}.$ (1)
Typically, the denoising network has a UNet structure [40], although alternative options exists [35]. For text-to-image generations, a text encoder $f(\cdot)$, usually frozen, transforms a prompt to a text embedding $c$. This embedding is subsequently input into the denoising network.
Text-to-image diffusion models can be broadly categorized to: pixel-based and latent-based models. Pixel-based models [42, 38, 1] directly model the pixel value, whereas latent-based models [39, 36] use a pre-trained Variational Autoencoder (VAE) to compress images into latent space, and let the diffusion models capture the distribution of latent variables. Latent diffusion models have demonstrated superiority in computational efficiency, making them a suitable candidate for low-cost applications such as on mobile devices. As a result, we primarily focus on the architecture design latent diffusion models, while the extension to pixel-based models is left for future exploration.
Refer to caption Refer to caption
UNet Conv. + Transformer
@@ 3.1 Preliminary: Diffusion Models
Figure 2: Model illustration of MobileDiffusion (MD). ($\mathsf{Conv.}$: Convolution. $\mathsf{SA}$: Self-attention (optional). $\mathsf{CA}$: Cross-attention. $\phi$: Non-linearity to calculate attention weights. $e$: Expansion factor in gated linear layer [44]. $\mathsf{DW}$: Depth-wise convolution.)
3.2 UNet Optimization
@@
As illustrated in the left panel of Figure 2, UNet architectures for text-to-image diffusion models commonly interleave transformer blocks and convolution blocks. In this section, we conduct a comprehensive investigation of the these two fundamental building blocks. Throughout the study, we control the training pipeline (e.g. data, optimizer) to study the effects of different architectures. We leverage two metrics, Fréchet Inception Distance (FID) and CLIP-ViT B/32 on MS-COCO 2017 30K, complemented by visual inspections to measure the quality of each model variation. We start with a baseline model, which adopts the UNet architecture of SD [39]. After training with our pipeline, the model obtains FID 12.45 and CLIP score 0.325 with the cfg weight set at 7.5, which is comparable with SD-1.5.
3.2.1 Transformer Blocks
In text-to-image diffusion models, a transformer block [51], as depicted in the right part of Figure 2, typically consists of a self-attention layer (SA) for modeling long-range dependencies among visual features, a cross-attention layer (CA) to capture interactions between text conditioning and visual features, and a feed-forward layer (FF) to post-process the output of attention layers. These transformer blocks hold a pivotal role in text-to-image diffusion models, serving as the primary components responsible for text comprehension. However, they also pose a significant efficiency challenge, given the computational expense of the attention operation. [15] introduced the UViT architecture, which places more transformer blocks at the bottleneck of the UNet. This design choice is motivated by that the attention computation is less resource-intensive at the bottleneck due to its lower dimensionality. At a high level, our architecture shares similarities with this approach. Nevertheless, we introduce several key modifications, elaborated below.
More transformers in the middle of Unet. Motivated by UViT[15], we adopt a strategic approach of relocating the parameters from high-resolution transformer blocks to lower resolutions while ensuring a consistent total parameter count. Maintaining this parameter count consistency can be achieved by reducing channel dimensions at lower resolutions. Our empirical observations indicate that this reduction does not adversely impact the quantitative metrics or the visual quality of the generated samples. Alternatively, to uphold the parameter count while accommodating more transformer blocks, one might consider reducing the width of the transformer. However, our investigations reveal that a substantial width of the transformer blocks is vital for optimal performance. Attempting to stack more transformer blocks with a smaller width proves detrimental, resulting in an evident degradation in visual quality characterized by poor object composition and intricate artifacts. These findings guide our design choice of incorporating more transformer blocks at lower resolutions while simultaneously reducing the number of channels at those resolutions.
With comparable model size, this modification improves run-time efficiency by 26% without quality drop.
Decouple SA from CA. In text-to-image diffusion models, self-attention plays a pivotal role in capturing long-range dependencies, although it comes with significant computational costs, especially at higher resolutions. For instance, at a resolution of $32\times 32$, the sequence length for self-attention is $1024$, resulting in an exceptionally high computational burden due to attention’s quadratic complexity with respect to sequence length. In prior work, such as [15, 36], an approach involves relocating entire transformer blocks from high to lower resolutions, essentially removing both self-attention and cross-attention layers at high resolution. However, we observe that naively doing so leads to a notable performance decline. Specifically, the FID increases to 17.87 from the 12.50 baseline, and the CLIP score decreases to 0.302 from 0.325.
@@ 3.2.1 Transformer Blocks
Upon further investigation, we discover that retaining cross-attention layers while discarding only the self-attention layers at high resolutions does not result in a performance drop. We conjecture that cross-attention is useful across different resolutions, as text guidance is crucial for both the global layout and local texture of the image. Notably, the computational cost of cross-attention at high resolution is significantly lower than that of self-attention, due to the smaller sequence length of text embeddings (e.g. 77 for SD). Consequently, removing only the self-attention layer yields a substantial efficiency boost. In light of these insights, we adopt a design that maintains performance while enhancing efficiency: entirely removing transformer blocks at the highest resolution ($64\times 64$); eliminating self-attention layers in the transformer blocks at $32\times 32$ resolution and the outer $16\times 16$ stack; retaining complete transformer blocks in the inner $16\times 16$ stack and the innermost bottleneck stack.
This strategy improves computational efficiency by 15% without compromising model quality.
@@
Share key-value projections. Within an attention layer, both the key and value are projected from the same input $x$, denoted as $K=x\cdot W_{K}$ and $V=x\cdot W_{V}$. Our experiments reveal that adopting a parameter-sharing scheme, specifically setting $W_{K}=W_{V}$ for self-attention layers, does not adversely affect model performance. Therefore, we choose to implement this parameter-sharing strategy, resulting in approximately a 5% reduction in parameter count.
@@
This is optimized for post-training mobile deployment.
@@
Finetune $\mathsf{softmax}$ into $\mathsf{relu}$. Given key $K$, query $Q$ and value $V$, attention calculates $\mathbf{x}=\phi(K^{\mathsf{T}}Q)V$, where the function $\phi(\cdot)$ is the $\mathsf{softmax}$. However, the $\mathsf{softmax}$ function, where $\mathsf{softmax}(\mathbf{x})=e^{\mathbf{x}}/\sum^{N}_{j=1}e^{x_{j}}$ and $\mathbf{x}=(x_{0},x_{1},...,x_{N})$, is computationally expensive due to non-efficient parallelization of exponentiation and summation across the sequence length. Conversely, point-wise activations like $\mathsf{relu}$ present a quicker alternative that does not rely on specific hardware optimizations, and it can serve as a viable substitute [53]. Consequently, we propose employing $\mathsf{relu}$ in our attention computation. An intriguing finding is that there is no need to train a $\mathsf{relu}$-attention model from scratch; instead, finetuning from a pre-trained $\mathsf{softmax}$-attention model proves sufficient, and this fine-tuning process can be accomplished quickly, for example, within 10,000 iterations. Visual ablation results for this modification are provided in Appendix. Interestingly, we even observe an improvement in FID compared to the $\mathsf{relu}$-attention model, improved from 12.50 to 12.31.
This is also optimized across different mobile platforms.
@@
Trim feed-forward layers. The initial UNet transformer block features a considerable expansion ratio of 8 in its feed-forward layers, leading to a dimensional surge from 1280 to 10240. However, such high dimensionality can be limiting in resource-constrained mobile applications. After thorough ablations, we found that trimming the expansion ratio to 6 results in nearly identical performance. With this adjustment, the FID score experiences only a slight increase from 12.31 to 12.58, while concurrently leading to 10% reduction in parameters count.
3.2.2 Convolution Blocks
Convolution blocks, in particular ResNet blocks [11], are deployed at each level of the UNet. While these blocks are instrumental for feature extraction and information flow, the associated computational costs, especially at high-resolution levels, can be substantial. In this context, we delve into several critical insights aimed at enhancing the efficiency of convolution blocks.

### v2

# MobileDiffusion: Instant Text-to-Image Generation on Mobile Devices

@@ Abstract
The deployment of large-scale text-to-image diffusion models on mobile devices is impeded by their substantial model size and high latency. In this paper, we present MobileDiffusion, an ultra-efficient text-to-image diffusion model obtained through extensive optimizations in both architecture and sampling techniques. We conduct a comprehensive examination of model architecture design to minimize model size and FLOPs, while preserving image generation quality. Additionally, we revisit the advanced sampling technique by diffusion-GAN, and make one-step sampling compatible to downstream applications trained on the base model. Empirical studies, conducted both quantitatively and qualitatively, demonstrate the effectiveness of our proposed technologies. With them, MobileDiffusion achieves instant text-to-image generation on mobile devices, establishing a new state of the art.
Figure 1: MobileDiffusion for (a) Text to image generation. (b) Canny edge to image, style LoRA and inpainting. Samples are all generated in one-step.
@@ 1 Introduction
Text-to-image diffusion models [40, 44, 47, 1, 45, 42] have exceptional capabilities in generating high-quality images conditioned on texts. These models serve as the foundation for a variety of applications, including image editing[15, 55], controlled generation [65, 39], personalized content generation [46, 13], video synthesis[4, 14], and low-level vision tasks[61, 26]. These large-scale models are essentially considered for necessarily running on servers with powerful neural compute units. Only a few work [27, 25] has barely touched running diffusion models on mobile devices, which remains an open challenge.
@@
We identify two primary factors causing the inefficiency of text-to-image diffusion models. Firstly, the complexity of the network architecture in text-to-image diffusion models involves a substantial number of parameters, often reaching to the billions, resulting in computationally expensive evaluations. Secondly, the inherent design of diffusion models requires iterative denoising to generate images, necessitating multiple evaluations of the model [17, 53]. These inefficiency challenges pose a significant barrier to deploy in resource-constrained environments, such as on mobile devices. As a result, despite the potential benefits, such as enhancing user experience, addressing emerging privacy concerns, and saving cost, this aspect remains relatively unexplored within the current literature.
In this paper, we aim to develop an ultra-efficient diffusion model suitable for a range of on-device applications, necessitating a design that is lightweight, rapid, and versatile enough to handle various downstream generative tasks, including in-painting, controllable generation, and personalized generation efficiently. We approach the identified challenges—architectural and sampling efficiency—through a divide-and-conquer strategy, addressing each separately.
The challenge of architectural efficiency in diffusion models has been rarely addressed in existing literature. Prior attempts have focused on eliminating redundant neural network blocks [27, 25] or reorganizing them to improve efficiency [18] but lack a comprehensive analysis of the model architecture’s components. Our research fills this gap by conducting an in-depth review of diffusion networks, leading to a carefully optimized model architecture. With fewer than 400 million parameters, our model achieves high-quality image generation, surpassing previous efforts in efficiency.
The second challenge, namely the sampling efficiency of diffusion models, has been an active research area. With advanced numerical solvers[33, 34, 2, 23, 51] or distillation techniques [48, 38, 52, 27, 35], the necessary number of network evaluations of sampling has been reduced significantly from several hundreds to less than 10. Despite these advances, the reduced evaluation steps can still pose challenges on mobile devices, where even minimal processing demands may prove too burdensome. Our attention, therefore, turns towards pioneering single-step diffusion generation methods [62, 49, 63]. These advancements have the potential to transform diffusion models into efficient one-step text-to-image generators. Yet, there remains an unmet need for these models to flexibly accommodate a variety of conditional generation tasks through the integration of additional modules trained separately. To close this gap, we explore the nuanced design space of diffusion-GAN hybrids [62, 49, 28, 60], a forefront category of one-step diffusion models gaining popularity. Specifically, our examination centers on the UFOGen framework [62], a recent innovation in this domain. Our work advances the discourse by conducting an in-depth analysis of UFOGen’s objective function and training methodologies, culminating in a refined approach for developing highly effective one-step diffusion GAN models. This endeavor aims to achieve a dual objective: to facilitate the creation of high-quality text-to-image synthesis and to ensure the seamless adaptability of the model for diverse downstream applications, thereby marking a significant stride towards realizing ultra-fast, versatile one-step diffusion models.
Our combined efforts in architectural design and sampling efficiency enable instant generation of high-quality $512\times 512$ images on mobile devices, notably achieving 0.2 second on an iPhone 15 Pro, which is about the average response time of human to visual stimulus. This achievement marks a significant leap over previous state-of-the-art in on-device text-to-image generation [27]. We introduce our model as MobileDiffusion, highlighting its potential as a foundational generative model for edge devices.
Our paper makes several key contributions to the field of on-device generative modeling, detailed as follows:
  1. 1.
    We present a comprehensive exploration of architectural efficiency in text-to-image diffusion models. Our work introduces a refined diffusion model architecture that is not only highly efficient and lightweight but also demonstrates rapid performance on mobile devices.
  2. 2.
    We study the design space of one-step diffusion-GAN models to present the best recipe for training such models, leading to high generative capabilities beyond text-to-image applications in one-single step. To our knowledge, our work pioneers the adaptation of a one-step diffusion model for a broad range of applications.
  3. 3.
    Synthesizing our efforts, we introduce MobileDiffusion, an ultra-efficient diffusion model framework specifically designed for on-device deployment.
@@ 2 Related works
Architecture Efficiency A limited number of prior works have discussed the architectural efficiency of diffusion models. Early diffusion models [10, 45, 47] adopted a UNet structure with transformer blocks. Recent works [3, 41] started to introduce vision transformers to diffusion. [18] proposed UViT, a UNet structure with a transformer backbone. With transformers establishing their advances, the differences of these methods lie in approaches to balance computation and cost, through UNet or patchify. [25] introduced an approach to distill larger diffusion models into smaller student models by selectively removing specific blocks from the teacher model. Meanwhile, [27] presented an efficient architecture search method. They trained a UNet with redundant blocks using robust training techniques [21, 64] and then pruned certain blocks based on metrics, resulting in an architecture suitable for distillation. In our work, we align with the insights of [18] regarding the transformer components in the UNet while introducing distinctive perspectives elaborated in Section 3. In contrast to approaches like [27, 25], our focus extends beyond mere block removal. Instead, we conduct a more nuanced analysis and modification of the architecture. Additionally, unlike [27, 25], we opt not to employ knowledge distillation from a larger model. Our observations indicate that training our model from scratch yields satisfactory results.
Sampling Efficiency Diverse strategies have emerged to reduce the required sampling steps of diffusion models, broadly falling into two categories. The first involves developing fast solvers to more efficiently solve the differential equation tied to the denoising process, thereby reducing the necessary discretization steps steps [33, 33, 23, 2, 12]. The second approach leverages knowledge distillation techniques to compress the sampling trajectory  [37, 48, 52, 27, 35]. Among them, consistency distillation [52, 35, 36] have shown promise in reducing the number of sampling steps to 4-8, while still facilitating downstream generative tasks, as highlighted by [59]. Nonetheless, attempts to further decrease the number of steps often result in diminished output quality. Recent innovations aim to revolutionize this area by enabling single-step generation [30, 62, 49, 63]. One important direction of research in one-step diffusion is diffusion-GAN hybrids [60], namely fine-tuning diffusion models with an adversarial objective [62, 49]. Despite these advances, the adaptation of one-step diffusion models for a broader array of downstream tasks remains an under-explored territory. A concurrent work [28] employs parameter-efficient adversarial fine-tuning. However, their resulting parameter-efficiently fine-tuned model does not facilitate one-step generation.
3 Designing a mobile-friendly diffusion architecture
In this section, we present our recipe for crafting highly efficient text-to-image diffusion models, which ultimately lead to sub-second generation on mobile devices. Following [45, 42], we adopt latent diffusion for its efficiency of learning text-guided generation in latent space. Our model design lies in an efficient diffusion network and a lightweight VAE decoder. We employ the CLIP-ViT/L14 [43] as the text encoder, which can run in a couple of milliseconds on-device.
@@ 3 Designing a mobile-friendly diffusion architecture
Figure 2: Model illustration of MobileDiffusion. ($\mathsf{Conv}$: Convolution. $\mathsf{SA}$: Self-attention (optional). $\mathsf{CA}$: Cross-attention. $\phi$: Non-linearity to calculate attention weights. $e$: Expansion factor in gated linear layer [50]. $\mathsf{DW}$: Depth-wise convolution.)
3.1 Diffusion network
Early diffusion models [10, 45, 47] adopted a UNet architecture with convolutions and attentions. Recent works [3, 41] introduced vision transformer (ViT) backbones to diffusion with a great potential in scaling up the models. We noticed that transformers are expensive for a large sequence length. It needs to be considered when designing models with budget constraints. SnapFusion [27] removes transformer blocks at the highest resolution in latent space. ViT [41] patchifies a spatial input to a sequence of tokens, with a reduced length.
@@ 3.1 Diffusion network
We follow the UViT [18] for designing our diffusion network with text guidance. As shown in Fig. 2, we leverage the UNet to reduce the sequence length and compute transformers efficiently. We made novel changes to the UViT architecture by conducting a comprehensive investigation of two fundamental building blocks: transformer and convolution. Throughout the study, we control the training pipeline (e.g. data, optimizer) to study the effects of different architectures. We optimize the network towards lower numbers of parameters and floating point operations (FLOPs), while preserving the quality. We leverage two metrics, Fréchet Inception Distance (FID) and CLIP-ViT B/32 on MS-COCO 2014 30K, complemented by visual inspections to monitor the quality of each model variation.
3.2 Optimize transformer
Scale up the backbone. Model optimization does not always indicate scaling down. By relocating model parameters, we can achieve more efficient compute. We noticed that transformers are efficient at low resolutions. Therefore, we scale up the transformer backbone in the UViT by moving more transformer layers to the bottleneck, while maintaining the total number of parameters. Additionally, the computation complexity of self-attention is $O(nd^{2}+n^{2}d)$, where $n$ is the sequence length, and $d$ is the channel dimension. At low resolutions (e.g. $16\times 16$), the channel dimension contributes more to the computation. Our empirical observations indicate that slightly reducing the channel dimension in the bottleneck does not adversely impact the quantitative metrics or the visual quality of the generated samples. On the other hand, attempting to stack more transformer blocks with a large reduction of channel dimension proves detrimental, resulting in an evident degradation in visual quality characterized by poor object composition and intricate artifacts. We found $1024$ is the best channel dimension with a good trade-off of efficiency and quality.
@@ 3.2 Optimize transformer
Decouple SA from CA. In text-to-image diffusion models, self-attention plays a pivotal role in capturing long-range dependencies, although it comes with significant computational costs at higher resolutions. For instance, at a resolution of $32\times 32$, the sequence length for self-attention is $1024$. In prior work [18, 42], self-attention and cross-attention layers are moved together when approaching efficient designs. Upon further investigation, we discover that retaining cross-attention layers while discarding only the self-attention layers at high resolutions does not result in a performance drop. We conjecture that cross-attention is useful across different resolutions, as text guidance is crucial for both the global layout and local texture of the image. Notably, the computational cost of cross-attention at high resolution is significantly lower than that of self-attention, due to the smaller sequence length of text embeddings (e.g. 77 for SD). Consequently, removing only the self-attention layer yields a substantial efficiency boost. In light of these insights, we adopt a design that maintains performance while enhancing efficiency: entirely removing transformer blocks at the highest resolution ($64\times 64$); eliminating self-attention layers in the transformer blocks at $32\times 32$ resolution and the outer $16\times 16$ stack; retaining complete transformer blocks in the inner $16\times 16$ stack and the innermost bottleneck stack.
@@
Share key-value projections. Within an attention layer, both the key and value are projected from the same input $x$, denoted as $K=x\cdot W_{K}$ and $V=x\cdot W_{V}$. Our experiments reveal that adopting a parameter-sharing scheme, specifically setting $W_{K}=W_{V}$ for self-attention layers, does not adversely affect model performance. Therefore, we choose to implement this parameter-sharing strategy, resulting in approximately a $5\%$ reduction in parameter count.
@@
Finetune $\mathsf{softmax}$ into $\mathsf{relu}$. Given key $K$, query $Q$ and value $V$, attention calculates $\mathbf{x}=\phi(K^{\mathsf{T}}Q)V$, where the function $\phi(\cdot)$ is the $\mathsf{softmax}$. However, the $\mathsf{softmax}$ function, where $\mathsf{softmax}(\mathbf{x})=e^{\mathbf{x}}/\sum^{N}_{j=1}e^{x_{j}}$ and $\mathbf{x}=(x_{0},x_{1},...,x_{N})$, is computationally expensive due to non-efficient parallelization of exponentiation and summation across the sequence length. Conversely, point-wise activations like $\mathsf{relu}$ present a quicker alternative that does not rely on specific hardware optimizations, and it can serve as a viable substitute [57]. Consequently, we propose employing $\mathsf{relu}$ in our attention computation. An intriguing finding is that there is no need to train a $\mathsf{relu}$-attention model from scratch; instead, finetuning from a pre-trained $\mathsf{softmax}$-attention model proves sufficient, and this fine-tuning process can be accomplished quickly, for example, within 10,000 iterations. Visual ablation results for this modification are provided in Appendix.
@@
Trim feed-forward layers. In feed-forward layers of a transformer, the expansion ratio is default to 4, which is further doubled with gated units [50]. This amplifies the parameter count significantly. For example, a channel dimension 1280 surges to 10240 after projection. Such high dimensionality can be limiting in resource-constrained mobile applications. After thorough ablations, we found that trimming the expansion ratio to 3 results in nearly identical performance. With this adjustment, the FID score experiences only a slight increase of 0.27, while concurrently leading to $10\%$ reduction in parameters count.
3.3 Optimize convolution